Given that the quadratic equation
(1-k)x^2+8x+1-k=0
has equal roots, find the possible values of k.

Answer by stanbon(75887)

Given that the quadratic equation
(1-k)x^2+8x+1-k=0
has equal roots, find the possible values of k.
----
If roots are equal, b^2-4ac = 0
----
a = 1-k
b = 8
c = 1-k
-----
b^2-4ac = 64 - 4(1-k)^2 = 64-4(1-2k+k^2) = -4k^2+8k+60
------
Solve:
-4k^2+8k+60 = 0
k^2 -2k - 15 = 0
(k-5)(k+3) = 0
Possible value of "k":
k = 5 or k = -3
========================
Cheers,
Stan H.
=============